OpenStudy (anonymous):

can someone help explain how to do this please? Write the equation y - 2 = 4(x + 5) in standard form. A. 4x + y = 7 B. 4x - y = 3 C. 4x - y = -22

OpenStudy (anonymous):

Standard form is ax+by=c where, 'a' is positive and all the coefficients are integers (not fractions or decimals) y-2=4(x+5) y-2=4x+20 y-22 = 4x (subtract 20 on both sides) -22 = 4x-y (subtract y on both sides) Answer: 4x-y = -22 (c)

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

um no i think i get that now thank you very much for explaining exactly how to do it. :)